OpenStudy (anonymous):

Write a function rule for the cost of movie tickets shown in the table below. Please show the steps to your solution. # of Tickets 2  3  5 Cost in $ 17.50  26.25  43.75

OpenStudy (mathmate):

Find the rate of change, m: (In this case, rate of change is the cost of a ticket) \(m=\frac{\Delta y}{\Delta x}\) Data points are: (2,17.50), (3,26.25),(5,43.75). Pick the first two ordered pairs and calculate the rate. Pick the last two ordered pairs and calculate the rate. If they are the same, then the cost of tickets is a linear function. You can verify your conclusion by drawing a graph. It should be a straight line through the three ordered pairs.

OpenStudy (mathmath333):

https://www.desmos.com/calculator/owwfupeu7t

OpenStudy (mathmath333):

so it should be a line \(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

i dont get it @mathmath333

OpenStudy (anonymous):

@ryanvarghese12790

OpenStudy (anonymous):

@perl ??????